kindly help me with these problems. please answer these and show the solutions. thank you
1. x/77x-121y - y/49x-77y
2. a/6a-9b - b/4a-6b

Answer by stanbon(75887)

1. x/77x-121y - y/49x-77y
= x/[11(7x-11y)] - y/[7(7x-11y)]
----------------------
lcm = 77(7x-11y)
----------
Write each fraction with lcm as its denominator:
=[x(7) - y(11)]/lcm
= 1/77
==============
2. a/6a-9b - b/4a-6b
= a/(3(2a-3b)) - b/(2(2a-3b))
-------------------------------------
lcm = 6(2a-3b)
-------------
= (2a-3b)/[6(2a-3b)]
= 1/6
===============
Cheers,
stan H.